1 Prove that a necessary condition that



w = f(z) = u(x, y) + iv(x, y)



be analytic in a region R is that the Cauchy-Riemann equations



∂u



∂y =



∂v



∂y ,



∂u



∂y = −



∂v



∂x



are satisfied in a region R where it is supposed that these partial derivatives are



continuous in R


1
Expert's answer
2021-11-21T14:25:06-0500

Let’s suppose that f(z) is differentiable in some region A and f(z) = f(x + iy) = u(x, y) + iv(x, y). We’ll compute f'(z) by approaching z first from the horizontal direction and then from the vertical direction. We’ll use the formula

"f'(z) = \displaystyle{{\lim}_{ ∆z→0}}\frac{ f(z + ∆z) − f(z)}{ ∆z }"


where  ∆z = ∆x + i∆y

Horizontal direction: ∆y = 0, ∆z = ∆x


"f'(z) = \displaystyle{{\lim}_{ ∆z→0}}\frac{ f(z + ∆z) − f(z)}{ ∆z } =f'(z) = \displaystyle{{\lim}_{ ∆z→0}}\frac{ f(x + ∆x+iy) − f(x+iy)}{ ∆x } ="


"= \displaystyle{{\lim}_{ ∆z→0}}\frac{ u(x + ∆x,y) +iv(x + ∆x,y)-(u(x,y)+iv(x,y))}{ ∆x } ="


"=\frac{\partial u}{\partial x}(x,y)+i\frac{\partial v}{\partial x}(x,y)"


Vertical direction: ∆x = 0, ∆z = i∆y


"f'(z) = \displaystyle{{\lim}_{ ∆z→0}}\frac{ f(z + ∆z) − f(z)}{ ∆z } ="


"=\displaystyle{{\lim}_{ ∆z→0}}\frac{ u(x,y + ∆y) +iv(x,y + ∆y)-(u(x,y)+iv(x,y))}{ i∆y } ="


"=\frac{1}{i}\frac{\partial u}{\partial y}(x,y)+\frac{\partial v}{\partial y}(x,y)=\frac{\partial v}{\partial y}(x,y)-i\frac{\partial u}{\partial y}(x,y)"


We have found two different representations of f'(z) in terms of the partials of u and v. If put them together we have the Cauchy-Riemann equations:


"f'(z) =\frac{\partial u}{\partial x}(x,y)+i\frac{\partial v}{\partial x}(x,y)=\frac{\partial v}{\partial y}(x,y)-i\frac{\partial u}{\partial y}(x,y)\implies"


"\implies \frac{\partial u}{\partial x}(x,y)=\frac{\partial v}{\partial y}(x,y)" and "-\frac{\partial u}{\partial y}(x,y)=\frac{\partial v}{\partial x}(x,y)"